A particle with a mass of 0.2 g and a charge of 16 μC is moving with a speed of 15 m/s towards a long wire with a current of 13 A. At a certain distance from the wire, it experiences a magnetic force of 8 ×10-9 N . What is this distance (in SI units)?
Answer #1

let,

mass of the particle, m=0.2g

charge q=16*10^-6 C

speed v=15 m/sec

current in wire i=13A

magnetic force F=8*10^-9 N


now,

F=q*v*B

F=q*v*(uo*i/2pi*r)

(8*10^-9)=(16*10^-6)*15*(4pi*10^-7*13)/(2pi*r)

====> r=0.078 m

distnace r=0.078 m or 7.8 cm